Can anyone help me with this problem?
x^2+3x+2=0
Any help would be great.

Found 2 solutions by ewatrrr, sudhanshu_kmr:
Answer by ewatrrr(24785)


Hi
x^2+3x+2=0
(x+2)(x+1)=0 Note:SUM of the outside terms(x) and the inside terms(2x) = 3x
(x+2)=0 x = -2
(x+1)=0 x = -1

Answer by sudhanshu_kmr(1152)


x^2+3x+2=0
=> x^2+2x+x+2=0
=> x(x+2) +1(x+2) = 0
=> (x+1)(x+2) = 0
so, x = -1 or -2